Use scale of 18mm=1km to calculate the total length of the trail run.
Hint use straight line distance between points and multiply with 1,2
1
Expert's answer
2020-05-27T11:15:14-0400

So what we have: 18 mm equal 1 km. It means that 1 mm contains 18 times less distance than 18 mm:


"s=1\text{ km}/18\text{ mm}=0.0556\text{ km/mm}."

Now assume that straight line distance between points on the map is 157 mm. To find how many kilometres it is in the real world, simply multiply the distance on the map in mm by scale factor "s":


"D_\text{real}=D_\text{on map}\cdot s=157\text{ mm}\cdot\frac{1\text{ km}}{18\text{ mm}}=8.72\text{ km}."